OpenStudy (anonymous):

if the volume of a sphere is 4,500 cubic units what is the radius of the sphere

OpenStudy (anonymous):

Chocies A.) 15 units B.) 60units C.)5units D.)30units

OpenStudy (anonymous):

4/3*Pi*r^3=4500 r^3=3375/Pi r=10.241 so A?